# Attention Layers Add Into Low-Dimensional Residual Subspaces

### v1

@@ Abstract
While transformer models are widely believed to operate in high-dimensional hidden spaces, we show that attention outputs are confined to a surprisingly low-dimensional subspace, where about 60% of the directions account for 99% of the variance–a phenomenon that is induced by the attention output projection matrix and consistently observed across diverse model families and datasets. Critically, we find this low-rank structure as a fundamental cause of the prevalent dead feature problem in sparse dictionary learning, where it creates a mismatch between randomly initialized features and the intrinsic geometry of the activation space. Building on this insight, we propose a subspace-constrained training method for sparse autoencoders (SAEs), initializing feature directions into the active subspace of activations. Our approach reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features, and can further extend to other sparse dictionary learning methods. Our findings provide both new insights into the geometry of attention and practical tools for improving sparse dictionary learning in large language models.
@@ 1 Introduction
Over the past years, mechanistic interpretability has shifted from a collection of proof-of-concept tools (Olsson et al., 2022; Wang et al., 2022; Meng et al., 2023; Gould et al., 2023) toward a fast-growing, scale-driven field (Templeton et al., 2024; Ameisen et al., 2025; Lindsey et al., 2025). This transformation is driven by a wave of sparse dictionary learning methods–such as sparse autoencoders (SAEs) and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025)–that once targeted small models but are now being pushed to larger architectures and wider model families. As these approaches scale in performance and model coverage, they provide increasingly complete and fine-grained explanations of neural network behavior (Lindsey et al., 2024a; Gao et al., 2024).
@@
In Section 4, we show that attention outputs exhibit a remarkably strong low-rank structure compared to multilayer perceptron (MLP) outputs and residual streams. Through singular value decomposition and intrinsic dimension analyses (Guth et al., 2023; Staats et al., 2025), we demonstrate that this phenomenon holds universally across layers, model families, and datasets, which is consistent with the universality hypothesis (Olah et al., 2020; Chughtai et al., 2023; Gurnee et al., 2024; Wang et al., 2025). We further trace the origin of this low-rank structure to the anisotropy of the output projection matrix $W^{O}$, which compresses the multi-head outputs into a lower-dimensional subspace.
In Section 5, we investigate how the low-rank nature of attention outputs interacts with SAE training. By evaluating the full suite of open-source SAEs from LlamaScope (He et al., 2024), we show that low intrinsic dimension strongly correlates with the number of dead features, suggesting a mismatch between random initialization and the low-dimensional geometry of the activations. To address this, we propose Active Subspace Initialization, which aligns SAE features with the active subspace of activations, substantially reducing dead features while improving reconstruction. Following Lindsey et al. (2024a) and Gao et al. (2024), we conduct scaling experiments, which further reveal that ASI achieves superior reconstruction across feature counts, and when combined with SparseAdam [^1] https://docs.pytorch.org/docs/stable/generated/torch.optim.SparseAdam.html, it achieves the best reconstruction in large scale and reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features trained on Llama-3.1-8B (Dubey et al., 2024).
@@ 2.1 Low-Rankness in Attention Mechanisms
Prior work has investigated various notions of “low-rankness” within attention mechanisms.: low-rank approximation of attention patterns (Wang et al., 2020; Tay et al., 2020; Raganato et al., 2020), low-rank parameterization for model compression (Noach and Goldberg, 2020; Hu et al., 2022), and the inherent low-rank bottleneck in single-head outputs (Bhojanapalli et al., 2020).
It is important to note that our perspective differs from these prior lines of work. We demonstrate that the multi-head self-attention outputs naturally exhibit a low-rank structure, revealing a distinct and under-explored phenomenon.
@@ 2.2 Linear Representation Hypothesis and Sparse Dictionary Learning Methods
The linear representation hypothesis posits that high-level concepts correspond to linear directions in representation space (Arora et al., 2018; Olah et al., 2020; Elhage et al., 2022; Park et al., 2024). Building on this view, a series of sparse dictionary learning methods have been proposed as interpretability tools, including sparse autoencoders and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025). These approaches aim to decompose activations into combinations of sparsely activated features, while adopting different strategies, depending on their interpretability objectives, to predict or approximate feature activations. Importantly, this hypothesis has been shown to hold across a wide range of model scales (Templeton et al., 2024; Lieberum et al., 2024; He et al., 2024), architectures (Wang et al., 2025), and modalities (Abdulaal et al., 2024).
@@ 3.1 Multi-Head Self-Attention and Notations
We consider a Transformer layer with multi-head self-attention (MHSA). Given input representations $X\in\mathbb{R}^{n\times d}$, where $n$ is the number of tokens and $d$ is the hidden size, each attention head $i$ computes:
@@
Let $Z=\text{Concat}[Z_{1},\dots,Z_{H}]\in\mathbb{R}^{n\times d}$ denote the concatenated output of all attention heads (Nanda and Bloom, 2022). The final attention output is computed by applying the output projection:
@@
This formulation shows that $O$ can be viewed as the sum of low-dimensional outputs from each head, projected into the residual stream space. $O$ serves as the attention block’s contribution to the residual stream.
@@ 3.2 TopK Sparse Autoencoders
In this work, we adopt the TopK sparse autoencoder (TopK SAE) variant introduced by Gao et al. (2024). Unlike standard SAEs that impose an $\ell_{1}$ penalty, TopK SAE enforces exact sparsity by keeping only the top-$k$ activations in the latent representation for each input. Formally, given an input vector $x\in\mathbb{R}^{d}$, the encoder produces
@@ 4.1 Quantifying Low-Rankness with Relative Singular Values
We consider activation matrix $\widetilde{A}\in\mathbb{R}^{n\times d}$, where each row corresponds to one token’s activation vector, and $n$ is the number of data points, while $d$ is the dimensionality of the activation space (e.g., the hidden size of the model). Unless otherwise specified, $\widetilde{A}$ refers to the mean-centered activations. We refer readers to Appendix B for more details of activations sources.
@@
for a given threshold $\tau\in(0,1)$. Since SVD yields the optimal low-rank approximation in terms of reconstruction error, this $k$ provides a principled way to assess how concentrated the activations are in a low-dimensional subspace [^3] We use 0.99 as the threshold in main text, results of some other thresholds are provided in Appendix D, with no influence to the conclusion. (Figure 1). We further compute the fraction of delta downstream loss recovered by different number of components. These metrics complement our central findings, offering a numerical characterization of low-rankness.
@@ 4.2 Empirical Evidence of Low-Rank Structure
Figure 2: Results for $1M$ samples from SlimPajama (Soboleva et al., 2023) fed into Llama-3.1-8B. (A) The attention outputs are the most low-rank, as indicated by the sharpest decay in singular values. (B) The fraction of loss recovered by different number of components, selected in descending order of singular values.
@@
Attention outputs exhibit significantly lower intrinsic dimensionality compared to other activation types, with an intrinsic dimension of around 60% of the total dimensionality. In contrast, MLP outputs and the residual streams show much higher intrinsic dimensions above 90% (Figure 1).
Rapid Spectral Decay in Attention Outputs
Attention outputs exhibit a sharply decaying singular value spectrum. This is quantitatively evidenced by the number of components retaining significant energy: only 1174 singular values exceed 5% of the maximum value in attention outputs, compared to 3276 for MLP outputs and 3943 for the residual streams (Figure 2(a)).
@@
Attention outputs demonstrates superior dimensional efficiency in downstream task performance. Compared to zero ablation, attention outputs requires only 25.0% of dimensions to recover 95% of the downstream loss. This contrasts sharply with MLP outputs and residual streams, which require 78.1% and 85.9% of dimensions respectively to achieve the same recovery level (Figure 2(b)).
More results of these metrics across different layers, models and datasets are shown in Appendix C.
@@ 4.3 Low-Rankness of Attention Outputs Results from the Output Projection Matrix
Recall that the attention output is computed as $O=ZW^{O}$, where $Z\in\mathbb{R}^{n\times d}$ is the concatenated output of all attention heads, and $W^{O}\in\mathbb{R}^{d\times d}$ is a learned linear projection. To understand how the variance in $O$ is shaped, we analyze the variance of $O$ along an arbitrary unit direction $\hat{e}\in\mathbb{R}^{d}$, given by:
@@
Figure 4: The number of dead features and the intrinsic dimension of each layer in Llama-3.1-8B, shows a surprising consistency: activations with lower intrinsic dimensions have more dead features.
@@ 5.1 Empirical Correlation Between Low-Rank Structure and Dead Features
To explore how low-rankness affects the interpretability of attention, we use the same framework and data as the original study to evaluate the LlamaScope SAEs (He et al., 2024) [^4] Another prominent open-source SAEs, GemmaScope (Lieberum et al., 2024), train their attention SAEs on Z rather than attention output., which provide a complete set of SAEs trained on attention output, MLP output, and residual stream. We find that the number of dead features is strongly related to intrinsic dimensions, as shown in Figure 4. This observation suggests that dead features may stem from a fundamental mismatch between the SAE’s randomly initialized weights and the geometry of the activation space.
@@ 5.2 Active Subspace Initialization for Sparse Autoencoders
To investigate this co-occurrence, we train SAEs on the attention output of Llama-3.1-8B at layer 15, initializing the SAE feature directions in the subspace spanned by the top $d_{\text{init}}$ singular vectors of the activations. As shown in Figure 5, we find that constraining the initialization to a lower-dimensional active subspace substantially decreases the number of dead features.
Based on this observation, we propose Active Subspace Initialization (ASI), a lightweight and generalizable strategy for scaling SAEs to high capacities. Let $d$ be the input dimension, $h$ the SAE hidden dimension, and $n$ the number of data points. Given activation matrices $\widetilde{A}\in\mathbb{R}^{n\times d}$, we compute the SVD:
$\widetilde{A}=U\Sigma V^{\top},\quad V\in\mathbb{R}^{d\times d}.$
We select the top $d_{\text{init}}$ right singular vectors to form the active subspace:
$V_{\text{active}}=V_{:,:d_{\text{init}}}\in\mathbb{R}^{d\times d_{\text{init}}}.$
We then initialize the SAE weights directly in this subspace:
$W_{d}\in\text{span}(V_{\text{active}}),\quad W_{e}=W_{d}^{\top}.$
Figure 5: Proportion of dead features and normalized MSE across different subspace dimensions. Random subspaces are used as the baseline, whereas only initialization with the active subspace yields improvement.
@@
Figure 6: At a fixed number of features ($n=32768$), Active Subspace Init achieves a better reconstruction-sparsity trade-off than Base TopK and AuxK. A similar trend is observed in its impact on Language Model Downstream Loss. Note: The improvement in downstream loss is less pronounced than that in reconstruction, likely because Active Subspace Init allocates more features to the active subspace, which implicitly enhances reconstruction quality.
Enhanced Sparsity-Reconstruction Frontier Without Additional Compute
It empirically outperforms TopK on the sparsity-reconstruction frontier. It achieves slightly better results compared to the auxiliary loss approach while introducing no additional computational overhead (Figure 6).
@@
Our approach demonstrates optimal scaling behavior across various model configurations. On SAEs with an expasion factor of 32 times or more, it can achieve the same or even better performance as the auxiliary loss approach with only half the number of parameters (Section 5.3).
Broad Architectural Applicability
The technique maintains applicability to diverse activation functions and architectural variants, as it operates directly on the intrinsic properties of activations. This generalizability is further explored in Section 5.4.

### v2

@@ Abstract
Transformer architectures, and their attention mechanisms in particular, form the foundation of modern large language models. While transformer models are widely believed to operate in high-dimensional hidden spaces, we show that attention outputs are confined to a surprisingly low-dimensional subspace, where about 60% of the directions account for 99% of the variance–a phenomenon that is consistently observed across diverse model families and datasets, and is induced by the attention output projection matrix. Critically, we find this low-rank structure as a key factor of the prevalent dead feature problem in sparse dictionary learning, where it creates a mismatch between randomly initialized features and the intrinsic geometry of the activation space. Building on this insight, we propose a subspace-constrained training method for sparse autoencoders (SAEs), initializing feature directions into the active subspace of activations. Our approach reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features, and can further extend to other sparse dictionary learning methods. Our findings provide both new insights into the geometry of attention and practical tools for improving sparse dictionary learning in large language models.
@@ 1 Introduction
Over the past years, mechanistic interpretability has shifted from a collection of proof-of-concept tools (Olsson et al., 2022; Wang et al., 2022; Meng et al., 2023; Gould et al., 2023) toward a fast-growing, scale-driven field (Ameisen et al., 2025; Lindsey et al., 2025). This transformation is driven by a wave of sparse dictionary learning methods–such as sparse autoencoders (SAEs) and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025)–that once targeted small models but are now being pushed to larger architectures and wider model families (Templeton et al., 2024; Gao et al., 2024; Hazra et al., 2025). As these approaches scale in performance and model coverage, they provide increasingly complete and fine-grained explanations of neural network behavior (Lindsey et al., 2024a; Gao et al., 2024).
@@
In Section 4, we show that attention outputs exhibit a remarkably strong low-rank structure compared to multilayer perceptron (MLP) outputs and residual streams. Through singular value decomposition and intrinsic dimensionality analyses (Guth et al., 2023; Staats et al., 2025), we demonstrate that this phenomenon holds universally across layers, datasets, and model families—Llama 3.1 (Dubey et al., 2024), Gemma 2 (Rivière et al., 2024), and Qwen 3 (Yang et al., 2025), which is consistent with the universality hypothesis (Olah et al., 2020; Chughtai et al., 2023; Gurnee et al., 2024; Wang et al., 2025). We further trace the origin of this low-rank structure to the anisotropy of the output projection matrix $W^{O}$, which compresses the multi-head outputs into a lower-dimensional subspace.
In Section 5, we investigate how the low-rank nature of attention outputs interacts with SAE training. By evaluating the full suite of open-source SAEs from LlamaScope (He et al., 2024), we show that low intrinsic dimensionality strongly correlates with the number of dead features, suggesting a mismatch between random initialization and the low-dimensional geometry of the activations. Drawing inspiration from Phan et al. (2025)’s principal component initialization for the first network layer, we propose Active Subspace Initialization, which aligns SAE features with the active subspace of activations, substantially reducing dead features while improving reconstruction. Following Lindsey et al. (2024a) and Gao et al. (2024), we conduct scaling experiments, which further reveal that ASI achieves superior reconstruction across feature counts, and when combined with SparseAdam [^1] https://docs.pytorch.org/docs/stable/generated/torch.optim.SparseAdam.html, it achieves the best reconstruction in large scale and reduces dead features from 87% to below 1% in Attention Output SAEs with 1M features trained on Llama-3.1-8B (Dubey et al., 2024).
@@ 2.1 Low-Rankness in Attention Mechanisms
Prior work has investigated various notions of “low-rankness” within attention mechanisms.: low-rank approximation of attention patterns (Wang et al., 2020; Tay et al., 2020; Raganato et al., 2020), low-rank parameterization for model compression (Noach & Goldberg, 2020; Hu et al., 2022), and the inherent low-rank bottleneck in single-head outputs (Bhojanapalli et al., 2020).
Different from these prior lines of work, we demonstrate that the multi-head self-attention outputs exhibit a low-rank structure, revealing a distinct and under-explored phenomenon.
@@ 2.2 Linear Representation Hypothesis and Sparse Dictionary Learning Methods
The linear representation hypothesis suggests that high-level concepts align with linear directions in representation space (Arora et al., 2018; Olah et al., 2020; Elhage et al., 2022; Park et al., 2024). Motivated by this view, various sparse dictionary learning methods have been developed for interpretability, including sparse autoencoders and their variants (Cunningham et al., 2023; Bricken et al., 2023b; Lindsey et al., 2024b), transcoders (Dunefsky et al., 2024; Ge et al., 2024), and low-rank sparse attention (He et al., 2025). These methods decompose activations into sparse feature combinations, while differing in their strategies for predicting or approximating feature activations. The hypothesis has been validated across diverse model scales (Templeton et al., 2024; Lieberum et al., 2024; He et al., 2024), architectures (Wang et al., 2025), and modalities (Abdulaal et al., 2024).
@@ 2.3 Dead Features in Sparse Dictionary Learning Methods
2.4 PCA-Inspired Network Initialization
A common practice applies PCA to input data for dimensionality reduction before network training (Hastie et al., 2009; Montavon et al., 2012; Jolliffe, 1986; Bishop & Nasrabadi, 2007). Recently, Phan et al. (2025) proposed PCsInit, which initializes the first layer weights of networks with top principal components of data—embedding the PCA transform directly into the network. This provides the model with a superior parameter set (Gu et al., 2025), boosting performance by construction.
@@ 3.1 Multi-Head Self-Attention and Notations
We consider a Transformer block with multi-head self-attention (MHSA) (Vaswani et al., 2017). Given input activations $X\in\mathbb{R}^{n\times d}$, where $n$ is the token count and $d$ is the model hidden size, each attention head $i$ computes:
@@
Let $Z=\text{Concat}[Z_{1},\dots,Z_{H}]\in\mathbb{R}^{n\times d}$ denote the concatenated output of all attention heads (Nanda & Bloom, 2022). The final attention output is computed by applying the output projection:
@@
This formulation shows that $O$ can be viewed as the sum of low-dimensional outputs from each head, projected into the residual stream space. $O$ is the attention block’s contribution to the residual stream.
@@ 3.2 TopK Sparse Autoencoders
In this work, we adopt the TopK sparse autoencoder (TopK SAE) introduced by Gao et al. (2024). Unlike standard SAEs that impose an $\ell_{1}$ penalty, TopK SAE enforces exact sparsity by keeping only the top-$k$ activations in the latent representation for each input. Formally, given an input vector $x\in\mathbb{R}^{d}$, the encoder produces
@@ 4.1 Quantifying Low-Rankness with Relative Singular Values
We consider activation matrix $\widetilde{A}\in\mathbb{R}^{n\times d}$, where each row corresponds to one token’s activation vector, and $n$ is the number of data points, while $d$ is the dimensionality of the activation space (e.g., the hidden size of the model). Unless otherwise specified, $\widetilde{A}$ refers to the mean-centered activations. We refer readers to Appendix D for more details of activations sources.
@@
for a given threshold $\tau\in(0,1)$. Since SVD yields the optimal low-rank approximation in terms of reconstruction error, this $k$ provides a principled way to assess how concentrated the activations are in a low-dimensional subspace [^3] We use 0.99 as the threshold in main text, results of some other thresholds are provided in Appendix F, with no influence to the conclusion. (Figure 1). We further compute the fraction of delta downstream loss recovered by different number of components (Section 4.2). These metrics complement our central findings, offering a numerical characterization of low-rankness.
For each type of activations, we collect 10 million activation vectors [^4] In rare cases, outlier activations inflate variance along some directions, making variance-based dimensionality estimates unreliable; to address this, we filter activations with a norm exceeding $5\sigma$ from the mean.. We empirically verify that this sample size is sufficient to produce stable and reproducible singular spectrum analysis.
@@ 4.2 Empirical Evidence of Low-Rank Structure
Figure 2: Results for 1010M tokens from SlimPajama (Soboleva et al., 2023) fed into Llama-3.1-8B. (a) The attention outputs are the most low-rank, as indicated by the sharpest decay in singular values. (b) Fraction of loss recovered using varying numbers of top singular components.
@@
Attention outputs has an intrinsic dimension of around 60% of the total dimensionality. In contrast, MLP outputs and the residual streams show much higher intrinsic dimensions above 90% (Figure 1).
Rapid Singular Spectral Decay in Attention Outputs
This is quantitatively evidenced by the number of components retaining significant energy: only 74.7% singular values exceed 1% of the maximum in attention outputs, versus 100.0% for MLP outputs and residual streams (Figure 2(a)).
@@
Compared to zero ablation, attention outputs requires only 39.1% of dimensions to recover over 99% of the downstream loss, versus 95.3% and 96.9% for MLP outputs and residual streams (Figure 2(b)).
More results of these metrics across different layers, models and datasets are shown in Appendix E.
@@ 4.3 Low-Rankness of Attention Outputs Results from the Output Projection Matrix
Recall that the attention output is computed as $O=ZW^{O}$, where $Z\in\mathbb{R}^{n\times d}$ is the concatenated output of all attention heads, and $W^{O}\in\mathbb{R}^{d\times d}$ is a learned linear projection. To understand how the variance in $O$ (singular value spectra of O) is shaped, we analyze the variance of $O$ along an arbitrary unit direction $\hat{e}\in\mathbb{R}^{d}$, given by:
@@ 5.1 Empirical Correlation Between Low-Rank Structure and Dead Features
Figure 4: The number of dead features (left) and the intrinsic dimension (right) of each activation in Llama-3.1-8B, shows a surprising consistency: activations with lower intrinsic dimensions have more dead features, corresponding to all layers of attention output and last two layers of MLP output.
To explore how low-rankness affects the interpretability of attention, we use the same framework and data as the original study to evaluate the LlamaScope SAEs (He et al., 2024), which provide a complete set of SAEs trained on attention output, MLP output, and residual stream  [^5] Another prominent open-source SAEs, GemmaScope (Lieberum et al., 2024), train their attention SAEs on Z rather than attention output.. We find that the number of dead features is strongly related to intrinsic dimensions, as shown in Figure 4. This observation suggests that dead features may stem from the low-rank geometry of the activation space.
@@ 5.2 Active Subspace Initialization for Sparse Autoencoders
Based on this observation, we propose Active Subspace Initialization (ASI), a lightweight and generalizable strategy for scaling SAEs to high capacities. Let $d$ denote the input dimension, $h$ the hidden dimension of the SAE, and $n$ the number of data points. Given activation matrices $\widetilde{A}\in\mathbb{R}^{n\times d}$, we compute the singular value decomposition (SVD):
$\widetilde{A}=U\Sigma V^{\top},\quad V\in\mathbb{R}^{d\times d},$
where $U\in\mathbb{R}^{n\times d}$ is the left singular vector matrix, $\Sigma\in\mathbb{R}^{d\times d}$ is the diagonal matrix of singular values, and $V\in\mathbb{R}^{d\times d}$ contains the right singular vectors.
We select the top $d_{\text{init}}$ right singular vectors to form the active subspace and then initialize the SAE weights directly in this subspace:
$V_{\text{active}}=V_{:,:d_{\text{init}}}\in\mathbb{R}^{d\times d_{\text{init}}},\quad W_{D}\in\text{span}(V_{\text{active}}),\quad W_{E}=W_{D}^{\top},$
where $W_{E}$ is the encoder weight matrix and $W_{D}$ is the decoder weight matrix. Intuitively, ASI aligns the initial SAE parameters with the active directions of the data, ensuring that SAEs start in a meaningful low-dimensional subspace. As $d_{\text{init}}$ decreases from the full space dimension [^6] Setting $D_{init}$ equal to the full space dimension is equivalent to not using Active Subspace Initialization. within a certain range, the number of dead features in the SAE rapidly drops, with a corresponding improvement in Delta LM loss [^7] This metric is defined as the difference between the original language model loss and the loss when the SAE is inserted at the corresponding position, evaluated over 1 million tokens. (Figure 5). Further ablation studies and implementation details are contained in Appendix B and Appendix I, respectively.
Figure 5: After using ASI, proportion of dead features (left), normalized MSE (mid) and Delta LM loss (right) across different subspace dimensions for activations with a full space dimension of 4096.
Figure 6: At a fixed number of features ($n=32768$), Active Subspace Init achieves a better reconstruction-sparsity trade-off than TopK and AuxK. A similar trend is observed in its impact on Delta LM Loss. Note: The improvement in downstream loss is less pronounced than that in reconstruction, likely because Active Subspace Init allocates more features to the active subspace, which implicitly enhances reconstruction quality.
@@
Reduced dead features and Enhanced Sparsity-Reconstruction Frontier Without Additional Compute
It achieves near-zero dead features and slightly superior results compared to the auxiliary loss approach (AuxK), at no additional computational cost of the same order. (Figure 6).
@@
Our approach demonstrates optimal scaling behavior across various SAE training methods. It outperforms TopK and AuxK in any evaluated scale, from 1616K to 11M features (Section 5.3).
General Applicability
The technique maintains applicability to diverse architectural variants and activation functions, as it operates directly on the intrinsic properties of activations. This generalizability is further explored in Section 5.4 and Appendix H.